### v3

# PaTAS: A Framework for Trust Propagation in Neural Networks Using Subjective Logic

Abstract
Trustworthiness has become a key requirement for the deployment of artificial intelligence systems in safety-critical applications. Conventional evaluation metrics, such as accuracy and precision, fail to appropriately capture uncertainty or the reliability of model predictions, particularly under adversarial or degraded conditions. This paper introduces the Parallel Trust Assessment System (PaTAS), a framework for modeling and propagating trust in neural networks using Subjective Logic (SL). PaTAS operates in parallel with standard neural computation through Trust Nodes and Trust Functions that propagate input, parameter, and activation trust across the network. The framework defines a Parameter Trust Update mechanism to refine parameter reliability during training and an Inference-Path Trust Assessment (IPTA) method to compute instance-specific trust at inference. Experiments on real-world and adversarial datasets demonstrate that PaTAS produces interpretable, symmetric, and convergent trust estimates that complement accuracy and expose reliability gaps in poisoned, biased, or uncertain data scenarios. The results show that PaTAS effectively distinguishes between benign and adversarial inputs and identifies cases where model confidence diverges from actual reliability. By enabling transparent and quantifiable trust reasoning within neural architectures, PaTAS provides a foundation for evaluating model reliability across the AI lifecycle.
I Introduction
Artificial Intelligence (AI) systems, particularly Neural Networks (NNs) [1] are being employed in critical sectors such as healthcare to interpret clinical images [2] or in autonomous driving to recognize physical elements in traffic images [3]. While highly performant, they often operate as black boxes that offer limited insight into the reliability of their outputs. This opacity becomes critical under adversarial, uncertain, or degraded input conditions. Conventional performance metrics such as accuracy or precision do not capture uncertainty and reliability, leaving decision-makers with misleading indicators of a model’s outputs quality [4]. Even confidence estimates rarely account for dataset bias, label noise, or adversarial corruption. For instance, a diagnostic model may report 95% confidence in a classification task while being evaluated on mislabeled data. These challenges underscore the need for reliability measures that go beyond predictive confidence and that consider the quality and provenance of inputs, the reliability of training data, and the stability of learned parameters.
As AI systems become integral to critical real-world applications, trustworthiness has emerged as a fundamental requirement for ensuring reliability, safety, and alignment with human values. According to the High-Level Expert Group on AI [5], a trustworthy system must be lawful, ethical, and robust. In technical terms, this translates into measurable properties such as accuracy, robustness, fairness, and explainability [6]. Embedding these properties across the AI pipeline, from data collection to deployment, is essential for preventing failures caused by poor data quality, bias, or unstable training. In this work, we adopt a property-based view of trustworthiness, where a property refers to a specific technical aspect such as accuracy or bias. We focus on quantifiable technical dimensions that can be modeled and propagated within neural networks. Specifically, we study how trustworthiness in the dataset, the query input, and the learned parameters (assessed for the same property) jointly determine the reliability of model predictions.
These issues become even more pronounced in transfer learning scenarios, where knowledge is reused across domains or tasks. Transfer learning introduces additional challenges related to both the transferability and reliability of knowledge. A recent study [7] highlights the need for frameworks that assess whether transferred knowledge remains robust, fair, and secure across varying data distributions. These findings emphasize the need for a uniform approach for evaluating and propagating trustworthiness in complex neural networks.
Despite this growing awareness, most models still provide limited mechanisms for representing confidence in their outputs. Neural networks are typically trained with the assumption of clean, unambiguous data labels and thus lack in the capability to generalize across previously unseen borderline cases. As a result, predicted probabilities tend to reflect similarity to seen patterns rather than genuine uncertainty, and they often assume clean, in-distribution inference inputs. These assumptions are rarely satisfied in real-world or adversarial settings, where data quality, distributional shifts, and parameter stability all influence prediction reliability. Consequently, existing approaches lack mechanisms to evaluate how uncertainty and trustworthiness in training inputs, activations, and parameters jointly affect model behavior. Even attribution-based methods (like SmoothGrad [8]) fail to capture how training input, intermediate, and parameter trustworthiness jointly influence the output reliability. This gap motivates the development of tools that can explicitly reason about trustworthiness across the entire model lifecycle.
Problem Statement
In summary, existing methods for trustworthiness and uncertainty estimation in neural networks face several key limitations:
  1. 1.
    Neglect of data provenance and quality: Most frameworks assume the training data are fully reliable.
  2. 2.
    Limited holistic propagation: Uncertainty quantification methods typically assess reliability only at the output layer.
  3. 3.
    Lack of interpretability: Few models yield trustworthiness estimates that are both faithful to the model’s reasoning and understandable to end users.
These limitations hinder reliable trustworthiness assessments, particularly in safety-critical or adversarial contexts.
Contributions
To address these challenges, we propose the Parallel trustworthiness Assessment System (PaTAS), a framework for modeling and propagating trustworthiness in NNs using Subjective Logic (SL). The main contributions are summarized as follows:
  1. 1.
    Parallel Trust Computation: We introduce Trust Nodes and Trust Functions that mirror neural computations, enabling principled trust propagation during training and inference through SL discounting and fusion.
  2. 2.
    Parameter Trust Update: We design an algorithm that determines trust in learned parameters using gradient values, input trust, and label trust, and aligning parameter reliability with the learning dynamics.
  3. 3.
    Inference-Path Trust Assessment (IPTA): We propose a context-aware trust function that leverages activation-path information to compute per-instance trust scores.
  4. 4.
    Empirical Validation: We evaluate PaTAS on real-world and adversarial datasets, demonstrating that it produces interpretable, symmetric, and convergent trust estimates that reflect both input quality and internal model behavior.
Structure of the Paper
The remainder of this paper is organized as follows: Section II introduces background on Subjective Logic and dataset trustworthiness assessment, Section III reviews related work, Section IV formalizes trust propagation in neural networks, Section V details the PaTAS architecture, Section VI presents experiments, Section VII discusses implications, and Section VIII concludes the paper and provides an outlook on future work.
II Background
Trust assessment in NNs requires a formalization that may represent ambiguity, inadequate evidence, and source reliability. Section II-A summarizes the principles of Subjective Logic (SL) and the trust opinion representation, which encodes trust, distrust, and uncertainty. Section II-B covers the primary SL reasoning operators and their application in Subjective Trust Networks. Finally, we explain in Section II-C how these concepts manifest as measures of dataset trustworthiness.
II-A Subjective Logic Fundamentals
Reasoning about trust in AI systems requires handling partial, conflicting, or missing evidence. Classical probability theory models aleatoric uncertainty but cannot represent missing knowledge (ignorance) or contradictory information, which often arises from noisy or biased data. Subjective Logic [9] extends Dempster–Shafer theory [10, 11] to capture these conditions by representing trust as SL opinions rather than probabilities. A subjective opinion expresses beliefs and uncertainty as separate components, thereby distinguishing between uncertainty due to lack of knowledge (epistemic uncertainty) and uncertainty inherent to the environment itself (aleatoric uncertainty).
A subjective opinion denoted by $\omega_{X}^{A}$ expresses the beliefs of an agent $A$ (e.g., a sensor, a human, or an external observer of a process) about states of a variable $X$ which takes its values from a domain $\mathbb{X}$ (i.e., a state space). A special case of a subjective opinion is a subjective binomial opinion where $card(\mathbb{X})=2$. For a binary variable $X\in\mathbb{X}=\{x,\overline{x}\}$, a binomial opinion is expressed as a quadruple:
$\omega_{X=x}=\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$
satisfying $b_{x}+d_{x}+u_{x}=1$, where $b_{x}$ denotes belief in $x$, $d_{x}$ disbelief in $x$ (belief in $\overline{x}$), $u_{x}$ the uncertainty mass, and $a_{x}$ the base rate (prior probability of $x$ in the absence of evidence). Its projected probability is defined as:
$P(x)=b_{x}+a_{x}u_{x}.$ (1)
This projection reduces the richer subjective opinion to an equivalent classical probability, enabling compatibility with standard probabilistic reasoning.
Binomial opinions can be derived from evidence through various quantification models. Let $r_{x}$ and $s_{x}$ represent the amount of positive and negative evidence. Positive evidence $r_{x}$ captures observations that increase confidence in the truth of $x$, while negative evidence $s_{x}$ captures observations that support its falsehood. Three common quantification approaches are:
  • Baseline-Prior Quantification:
    $b_{x}$ $=\frac{r_{x}}{W+r_{x}+s_{x}},\quad d_{x}=\frac{s_{x}}{W+r_{x}+s_{x}},$
    $u_{x}$ $=\frac{W}{W+r_{x}+s_{x}}$ (2)
    where weight $W>0$ guarantees residual uncertainty.
  • Evidence-Weighted Quantification [12]:
    $b_{x}$ $=\frac{r_{x}}{w_{x}+r_{x}+s_{x}},\quad d_{x}=\frac{s_{x}}{w_{x}+r_{x}+s_{x}},$
    $u_{x}$ $=\frac{w_{x}}{w_{x}+r_{x}+s_{x}}$ (3)
    where the uncertainty is scaled by $w_{x}$.
  • Constant-Uncertainty Quantification [12]:
    $u_{x}$ $=U,\quad\gamma=\frac{1-U}{r_{x}+s_{x}}$
    $b_{x}$ $=\gamma\cdot r_{x},\quad d_{x}=\gamma\cdot s_{x}$ (4)
    where a fixed uncertainty distributes the remaining mass.
A subjective opinion is meaningful only within a specific context or property under evaluation (e.g., accuracy, bias, or other trust-related aspects). In this work, trust is represented as a subjective binomial opinion $(t,d,u)$, where $t$ denotes trust (belief), $d$ distrust (disbelief), and $u$ uncertainty. These fundamentals define how trust is represented and interpreted as subjective opinions. To make them operational, Subjective Logic provides operators for combining, revising, and discounting opinions, which we introduce next.
II-B Subjective Logic Operators
SL provides key reasoning operators for combining and propagating opinions, including trust discounting, fusion, and inferential operators [13, 14, 9].
Definition 1 (Fusion [13]).
Let $A$ be an agent forming an opinion about a proposition $X=x$ based on two information sources, $P$ and $Q$. The fused opinion is defined as:
$\omega^{A}_{X=x}=\omega^{P\&Q}_{X=x}=\omega^{P}_{X=x}\odot\omega^{Q}_{X=x}.$ (5)
The specific fusion operator depends on the relationship between the sources. SL defines several variants such as consensus, averaging, weighting, and cumulative fusion, each representing a distinct way of aggregating evidence.
Example. Suppose two temperature sensors estimate whether the room temperature exceeds $25^{\circ}$C. If both sensors are of the same type and installed in the same place, their evidence is correlated and averaging or weighted fusion is appropriate. If they have different measurement strategies (e.g., infrared and contact-based), cumulative fusion is more suitable. The information from each independent sensor is treated as an additional, non-redundant contribution so that the combined evidence grows with each agreeing source.
Definition 2 (Trust Discounting [14]).
Let $A$ have a referral trust $\omega^{A}_{B}$ in another agent $B$, who holds an opinion $\omega^{B}_{X}$ on variable $X$. The trust-discounted opinion of $A$ derived from $B$’s opinion is:
$\omega^{[A;B]}_{X=x}=\omega^{A}_{B}\otimes\omega^{B}_{X=x}.$
Referral trust is domain-specific and expresses how much $A$ relies on $B$ regarding $X$.
Example. Consider a monitoring system where agent $A$ receives readings from a temperature sensor $B$. The sensor usually works well but is known to drift at times, so $A$ does not fully trust it. When $B$ reports that the temperature is above a safety threshold, $A$ discounts this opinion by reducing its strength and increasing uncertainty. This illustrates the principle of trust discounting: evidence from a partially reliable source is treated cautiously.
Definition 3 (Inferential Operators [15]).
Inferential operators generalize Bayesian reasoning by enabling opinion propagation through conditional relationships. Let $A$ be an agent reasoning about a variable $X$ and its potential implications for another variable $Y$. Suppose $A$ holds opinions $\omega^{A}_{Y\mid X}$ on the relationship $X\Rightarrow Y$. The main operators are:
  • Deduction: derive $\omega^{A}_{Y}$ from $\omega^{A}_{X}$ using $\omega^{A}_{Y\mid X}$.
  • Abduction: derive $\omega^{A}_{X}$ from $\omega^{A}_{Y}$ using $\omega^{A}_{Y\mid X}$.
Example. For the relation “if it rains ($X$), then Bob carries an umbrella (YY),” agent $A$ holds a conditional opinion $\omega^{A}_{Y\mid X}=(b,d,u)$. Given an opinion on rain, deductive inference yields an opinion about umbrella use; conversely, abduction infers the likelihood of rain from umbrella observations.
A summary of all operators used in this work and their symbols appears in Appendix A.
Subjective Trust Networks
A Subjective Trust Network (STN) [16, 17] models trust relationships as subjective opinions propagated through referral chains using trust discounting and fusion. Such extensions of SL have been applied to trust propagation, opinion dynamics, and source reliability analysis [18, 19, 20, 21].
II-C Trustworthiness in the dataset
Focusing back on machine learning, the quality and structure of the training dataset are essential for determining the performance, robustness, and fairness of machine learning models. Common issues such as sampling biases, mislabeled instances, or lack of diversity in the data can degrade learned representations and hinder generalization, thereby reducing the reliability of the model’s outputs [22]. A recent study on dataset quality shows that even small proportions of mislabeled samples can substantially shift model rankings; for instance, on CIFAR-10, VGG11 trained on clean data can outperform VGG19 once the fraction of erroneous labels reaches about 5%, illustrating how sensitive benchmark conclusions are to label errors [23]. Thus, evaluating the trustworthiness of training data is a critical step in assessing the trustworthiness of an AI system.
Subjective Logic has been applied effectively to model dataset trustworthiness [12], treating the dataset as a collection of samples. Each sample consists of an input vector and its corresponding label. Thus, dataset trustworthiness can be assessed at various granularities depending on the sub-property of interest:
  • Dataset level: Captures global explanations of misbehavior such as class imbalance or sampling bias. These factors affect the overall distribution and may harm generalization.
  • Instance level: Captures local explanations of anomalies like mislabeled or corrupted data points. Individual instances may be unreliable due to noise, annotation errors, or improper data collection.
  • Input feature level: Captures fine-grained variations within the input vector. For instance, in a data poisoning scenario, an adversarial patch affecting only one specific pixel may render part of an image untrustworthy [24]. Similarly, when input feature values are sourced from heterogeneous systems, some features may be inherently more reliable than others, leading to variable trust across features.
III Related Work
Quantifying trust and uncertainty in neural networks has become a central research topic, particularly in safety-critical domains. Incorrect yet confident predictions can have severe consequences. For example, data poisoning occurs when both the training and test datasets contain systematically mislabeled samples. However, the model may obtain good accuracy, precision, and recall on the poisoned test set. This observation creates the illusion of dependability despite having learned harmful or erroneous patterns. Standard metrics therefore fail to capture subtleties of model trustworthiness. As a result, diverse frameworks have been proposed to model and propagate uncertainty and trust, yet significant limitations remain as we will point out next.
III-A Uncertainty Quantification in Neural Networks
Uncertainty quantification (UQ) aims to estimate predictive reliability by modeling uncertainty at different levels, typically epistemic (model-based) and aleatoric (data-based) [25]. Foundational methods include Bayesian neural networks [26, 27, 28], Monte Carlo dropout [29], and ensembles [30]. Dropout provides an efficient Bayesian approximation [29], while Bayes by Backprop [27] learns weight distributions that enhance generalization and exploration in reinforcement learning. Extensions with latent variables explicitly decompose predictive uncertainty into epistemic and aleatoric components [28], improving decision-making in active and reinforcement learning through risk-sensitive criteria.
Despite these advances, estimating and calibrating both forms of uncertainty in complex models remains challenging [31]. Studies report frequent over- or under-confidence, with uncertainty estimates often degrading under real-world conditions [32, 33, 34]. In medical AI, uncalibrated confidence has been linked to critical misjudgments [2, 32]. Even well-performing models can produce unreliable confidence scores under dataset shifts [34]. Consequently, researchers have explored post-hoc calibration, such as temperature scaling [35], which adjusts output probabilities to better align predicted and observed frequencies. However, these approaches operate only at the output layer and assume clean data, which can still yield misleading confidence when inputs or training data are corrupted.
III-B Subjective Logic Approaches to Trust
Subjective Logic (SL) provides a probabilistic framework for modeling belief, disbelief, and uncertainty, offering a structured approach to trust reasoning. Evidential deep learning [36] applies SL principles by representing class predictions as subjective opinions parameterized through a Dirichlet distribution. The model jointly predicts outcomes and quantifies confidence, distinguishing between low-confidence predictions and high-uncertainty regions such as out-of-distribution inputs. Although effective, this method assumes clean data and lacks input-level trust assessment, while PaTAS directly models such factors.
Other SL-based methods focus on interpretable trust quantification. A calibration-based approach [37] clusters model outputs into subjective opinions to derive per-prediction trust scores without accessing internal parameters. While simple and model-agnostic, it again assumes trustworthy datasets and neglects input evidence. The DeepTrust framework by Cheng et al. [38] instead adopts a white-box perspective, integrating dataset evidence during training to assess global model trustworthiness. Although holistic, DeepTrust’s use of SL fusion and multiplication operators lacks algebraic consistency. In particular, it maps neural-network addition to Subjective Logic fusion and neural-network multiplication to SL opinion multiplication, even though these two SL operators belong to different algebraic domains, since fusion operates over agents’ opinions whereas multiplication operates over variables, without clear explanation. Moreover, its formulation of trust backpropagation is specified only at a single-layer level, making its theoretical extension to deeper architectures unclear despite empirical evaluations on complex networks.
PaTAS addresses these limitations by introducing a coherent and well motivated, layer-wise propagation mechanism compatible with deep architectures.
III-C Trust and Uncertainty Propagation
Recent work on uncertainty propagation seeks to improve both accuracy and computational efficiency. Mae et al. [39] proposed a sampling-free conversion of dropout-trained networks into Bayesian models using variance propagation. Monchot et al. [40] employed Gaussian Mixture Models and a Split-and-Merge algorithm with a Wasserstein criterion to propagate input uncertainty without assuming Gaussianity, achieving convergence guarantees at low cost. Astudillo and Net [41] extended these ideas to multi-layer perceptrons for speech recognition, showing that observation uncertainty enhances robustness even in hybrid MLP-HMM systems.
Beyond neural architectures, Ziegler and Lausen [42] proposed the Appleseed model for trust propagation in social networks using dynamic spreading activation. Though not originally intended for neural systems, it demonstrates the value of viewing trust as a structural, context-dependent quantity, an idea further developed in PaTAS.
Existing approaches demonstrate growing interest in uncertainty and trust modeling to assess trust in neural networks, yet they often neglect the joint influence of input quality, data reliability, and model parameters on prediction trust. The trustworthiness of an output cannot be viewed as a fixed property of the model alone but must depend on the corresponding input and its propagation through the network. PaTAS addresses these gaps by modeling trust as a dynamic property distributed across inputs, parameters, and activations, enabling consistent and interpretable trust propagation that reflects both data quality and network structure.
IV Proposed Method
In our framework, trust reasoning begins at the feature level, where each individual input component is assigned a trust opinion. This design choice provides greater flexibility and fine-grained control, enabling the framework to reflect nuanced variations in input reliability across different operational contexts. For example, certain regions or even individual pixels of an input image could be individually assigned different trust opinions, derived from measurable indicators such as noise levels, blur estimates, or confidence measures produced by the imaging pipeline. These initial trust opinions are then injected into the PaTAS, which propagates them through its network. This propagation mechanism ensures that variations in input trust are explicitly carried through to the model’s outputs, enhancing the interpretability and transparency of AI decisions.
IV-A Foundations of Trust-Aware Neural Inference for PaTAS
This section formalizes the structure and behavior of a trust propagation framework that mirrors standard neural network computations.
Given a neural network represented by
$\Theta=\left(\mathbf{W},\mathbf{b},f\left(\cdot\right)\right)$
where:
  • The parameter $\mathbf{W}=\{\mathbf{W}^{(l)}\}$ is a list of weight matrices $\mathbf{W}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ for each layer $l$, with $n_{l}$ the number of neurons in layer $l$.
  • $\mathbf{b}=\{\mathbf{b}^{(l)}\}$ where $\mathbf{b}^{(l)}\in\mathbb{R}^{n_{l}}$ is the bias vector for layer $l$.
  • $f(\cdot)=\{f^{l}(\cdot)\}$ where $f^{l}(\cdot)$ is the activation function for layer $l$ which may vary across each neuron of the layer.
The network output $\mathbf{y\prime}$ for an input $x$ is computed from the standard feedforward equation:
$\mathbf{y\prime}=f_{\Theta}(x)$ (6)
$\mathbf{y\prime}=f_{L}\left(\mathbf{W}^{(L)}\left(f_{L-1}\left(\dots f_{2}\left(\mathbf{W}^{(2)}f_{1}\left(\mathbf{W}^{(1)}\mathbf{x}+\mathbf{b}^{(1)}\right)+\mathbf{b}^{(2)}\right)\dots\right)\right)+\mathbf{b}^{(L)}\right)$
Given a training dataset $\mathcal{D}=\left\{(\mathbf{x}_{i},\mathbf{y}_{i})\right\}_{i=1}^{N}$, used to train the neural network, and a trust assessment function $T(.)$ evaluating the trustworthiness of each features and labels data ($\mathbf{x}_{i}$ and $\mathbf{y}_{i}$), our objective is to compute a corresponding trust opinion on the network output $\mathbf{y\prime}=f_{\Theta}(\mathbf{x})$. To formalize this, we introduce the notion of a Parallel Function.
Definition 4 (Parallel Function).
The Parallel Function of a Neural Network with parameters $\Theta$, denoted $\mathit{Pf}_{\Theta}$, is a function that mirrors the structure of the network’s feedforward computation $f_{\Theta}$ to propagate trust assessments from input to output. Given a trust evaluation $T(\mathbf{x})$ over an input $\mathbf{x}$, $\mathit{Pf}_{\Theta}(T(\mathbf{x}))$ returns a trust opinion on the network’s output $\mathbf{y\prime}=f_{\Theta}(\mathbf{x})$. This opinion reflects the trust assigned to the prediction based on trust in the input, and also taking into account the architecture and how the parameters of $f_{\Theta}$ were learned.
To effectively construct $\mathit{Pf}_{\Theta}$, we must first understand how the underlying neural network is built.
The standard training objective is defined by the following optimization:
$\mathbf{W},\mathbf{b}=\arg\min_{\mathbf{W},\mathbf{b}}\sum_{i=1}^{N}\mathcal{L}\left(\mathbf{y\prime}_{i},\mathbf{y}_{i}\right)$ (7)
where:
  • $\mathbf{y\prime}_{i}=f_{\theta}(\mathbf{x}_{i})$ is the model output,
  • $\mathbf{y}_{i}$ is the true label,
  • $\mathcal{L}$ is a loss function, such as mean squared error or cross-entropy,
  • NN is the total number of samples in the dataset $\mathcal{D}$.
This optimization seeks to find the set of weights and biases that minimize the loss function across the entire training set $\mathcal{D}$, effectively improving the network’s ability to make accurate predictions.
While the training process optimizes model accuracy, it does not account for how trust in the input data influences trust in the output predictions. Therefore, accuracy alone is not enough to assess trustworthiness of a model. We therefore turn to Subjective Logic as a formal calculus for modeling and propagating trust through neural networks. As a first step in constructing the parallel function $\mathit{Pf}_{\Theta}$, we consider a simple perceptron model to analyze how input trust opinions can be propagated to the output through the network’s structure and parameters. This forms the basis for progressively building the complete formulation of $\mathit{Pf}_{\Theta}$.
IV-B Perceptron Case: SL Formulation
Figure 1: STN used to assess the output $y\prime=f(x)$ of a perceptron $\Theta$
Assume that an observer $A$ wants to form a trust opinion $\omega_{y\prime}^{A}$ on the output $y\prime$ of a perceptron. Since $A$ does not directly observe or interact with the internal process used to produce $y\prime$, their opinion must be inferred indirectly through the output neuron (or output neurons of the network in the general case), denoted $N_{O}$. Specifically, the observer relies on the trust opinion $\omega_{y\prime}^{N_{O}}$ formed by the output neuron, and holds a referral trust $\omega_{N_{O}}^{A}$, which expresses how much $A$ trusts $N_{O}$ (or how much $A$ trusts the process used by $N_{O}$) for providing good trust opinion on $y\prime$. Fig. 1 illustrates the corresponding STN.
The goal now is to compute $\omega_{y\prime}^{N_{O}}$. For that end, we state that trust in the output is impacted by trust in the input, and the trust in the perceptron itself. The trust in the perceptron is, in turn, influenced by the perceptron design and the trust in the training dataset.
Let $f_{\Theta}$ be the perceptron inference function, and let $x$ be an input with an associated trust opinion $T_{x}$ [^1] We use $T_{x}$ notation instead of $\omega_{x}$ to emphasize that, as input to the framework, only the opinion on the variable is required and no specific agent needs to be represented.. Given the output $y\prime=f_{\Theta}(x)$, our goal is to construct $\mathit{Pf}_{\Theta}(T_{x})$, the corresponding trust opinion on $y\prime$.
To explore this construction, we consider a simple perceptron model that estimates the cost of renting an apartment:
$y\prime=10\times s+100\times n_{r}\,$ (8)
where $y\prime$ is the size of the apartment and $n_{r}$ is the number of rooms.
Based on this model, we introduce two initial sub-problems to illustrate how trust propagates through the computation:
  • Problem 1: How does trust in $n_{r}$ and $s$ propagate to the output $y\prime$.
  • Problem 2: Assume that we have trust $T_{\theta_{1}}$ in $\theta_{1}=10$ and $T_{\theta_{2}}$ in $\theta_{2}=100$. Here, trust again depends on the property of interest. For example in a bias context, it reflects the extent to which the weights were trained to capture the true influence of apartment size and number of rooms on the final price, without introducing systematic bias. The central question, then, is how such parameter-trust assessments refine the solution of Problem 1. In other words, how does trust in the parameters influence the propagation of input trust to the output $y\prime$?
The question of how to calculate trust in the parameters is addressed in Section V.
IV-B1 Solution to Problem 1 - Trust Propagation from Input to Output
Figure 2: Trust-propagation based solely on input-feature trust
Objective
Determine how trust in the individual input features ($s$ and $n_{r}$) propagates through a simple perceptron to produce a trust assessment on the output variable $y\prime$, representing the predicted apartment cost.
Assumptions
For now, we assume that the model parameters $\theta_{1}=10$ and $\theta_{2}=100$ are fully trusted and input feature trust opinions ($T_{s}$ and $T_{n_{r}}$) are available.
The model specified in Eq. 8 is equivalent to the neural network depicted in Fig. 2(a). This network employs a linear transformation without an activation function. The input vector is:
$x=\begin{pmatrix}s\\ n_{r}\end{pmatrix}$
Let $\omega_{x}^{s}$ be the trust opinion on $x$ based on $s$ as evidence and $\omega_{x}^{n_{r}}$ the trust opinion on $x$ based on $n_{r}$ as evidence. Since the output neuron $N_{3}$ computes Eq. 8, we associate this computation with two agents: $10N_{1}$ from the context of computing $10\cdot s$ and $100N_{2}$ from the context of computing $100\cdot n_{r}$. The trust opinion of the output neuron $N_{3}$ on $x$ is then:
$\omega_{x}^{N_{3}}=\mathit{fusion}(\omega_{x}^{10N_{1}},\omega_{x}^{100N_{2}})$
The choice of the fusion operator (that we will later denote by $\oplus$ ) depends on the semantics of $s$ and $n_{r}$ (see Definition 1). In this example, since $s$ and $p$ represent independent evidence, the fusion operator to use is cumulative fusion.
Assuming full trust in the parameters $\theta_{1}=10$ and $\theta_{2}=100$, we have:
$\omega_{x}^{10N_{1}}=\omega_{x}^{N_{1}}=\omega_{x}^{s}=T_{s}\,\text{ (Trust in the feature $s$ of $x$)}$ (9)
$\omega_{x}^{100N_{2}}=\omega_{x}^{N_{2}}=\omega_{x}^{n_{r}}=T_{n_{r}}\,\text{ (Trust in the feature $n_{r}$ of $x$)}$ (10)
Since the output value is deterministically related to the input via
$y^{\prime}=f(x),$
the trust opinion on $y^{\prime}$ is the same as the trust opinion already computed for $x$. Thus,
$\omega_{y\prime}^{N_{3}}$ $=\omega_{x}^{N_{3}}=\omega_{x}^{10N_{1}}\oplus\omega_{x}^{100N_{2}}$
$=\omega_{x}^{s}\oplus\omega_{x}^{n_{r}}=T_{s}\oplus T_{n_{r}}$ (11)
In summary, if we define $T_{x}=\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix}$, then:
$\mathit{Pf}_{\Theta}(T_{x})=\mathit{Pf}_{\Theta}(\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix})=T_{s}\oplus T_{n_{r}}$
IV-B2 Solution for Problem 2 - Impact of Parameter Input Trust on Output Trust
Objective
Analyze how trust in the model parameters $\theta_{1}=10$ and $\theta_{2}=100$ affects the resulting trust assessment on the output $y\prime$, given trust in the inputs.
Assumption
Trust Opinions $T_{\theta_{1}}$ and $T_{\theta_{2}}$ are assigned to the parameters. The approach used to calculate these parameter-trust values will be discussed in Section V.
In problem 2, the model is expressed as:
$y\prime=\theta_{1}\times s+\theta_{2}\times n_{r}$ (12)
and we assume trust parameters $T_{\theta_{1}}$ and $T_{\theta_{2}}$ respectively in $\theta_{1}$ and $\theta_{2}$ (we’ll see in details in Section V how to calculate these trust parameters). Unlike the previous problem, where we fully trusted $\theta_{1}$ and $\theta_{2}$, here we do not fully trust these parameters, and we must account for their trust assessments. For the network to incorporate the trust in $\theta_{1}$ and $\theta_{2}$, we adjust the trust opinions on the features accordingly. The trust opinions are now refined.
Figure 3: Trust-propagation Trust Network including parameter trust.
As depicted in Fig. 3, we model this as a small subjective network. therefore:
$\omega_{x}^{\theta_{1}N_{1}}=T_{\theta_{1}}\otimes\omega_{x}^{N_{1}}\;\text{and}\;\omega_{x}^{\theta_{1}N_{2}}=T_{\theta_{2}}\otimes\omega_{x}^{N_{2}}$ (13)
where $\otimes$ is a trust discounting operator. This results is consistent with the solution for problem 1 as for fully trusted $T_{\theta}=(1,0,0)$, we have $\omega_{x}^{\theta N}=T_{\theta}\otimes\omega_{x}^{N}=\omega_{x}^{N}$ (for any neuron $N$).
Finally, the resulting output trust $\omega_{y\prime}^{N_{3}}$ is calculated by the fusion of these adjusted trust opinions:
$\mathit{Pf}_{\Theta}(T_{x})=\mathit{Pf}_{\Theta}(\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix})=(T_{\theta_{1}}\otimes T_{s})\oplus(T_{\theta_{2}}\otimes T_{n_{r}})$ (14)
Thus, we adjust the trust in the network output based on the trust in both the parameters and the features, ensuring that the trust propagation takes into account the trust in the parameters.
IV-C Trust Nodes and Trust Functions
(a) Structure of a Perceptron
(b) Mirrored Trust Structure
(c) Perceptron Operation
(d) Trust Function Operation
Figure 4: Illustration of the Trust Node structure and computation
We now need to extend our discussion of $\mathit{Pf}_{\Theta}$ from a single perceptron to larger neural networks. In order to formalize the construction of $\mathit{Pf}_{\Theta}$, we introduce two fundamental concepts: the Trust Node (Fig. 4(b)) and the Trust Function (Fig. 4(d)). These components provide the basic mechanisms for propagating trust through the structure of a neural network.
Definition 5 (Trust Node).
A Trust Node is an abstract computational unit associated with a neuron in a neural network. It receives trust opinions on the neuron’s inputs and produces a trust opinion on the neuron’s output. The structure of a Trust Node mirrors that of its corresponding neuron, but its computation is defined over trust opinions using operators such as discounting and fusion.
Definition 6 (Trust Function).
A Trust Function models the transformation of trust through a Trust Node. It defines how trust opinions on the inputs of a neuron are combined to produce a trust opinion on the output. For a neuron that computes
$z^{(l)}=\sum\theta^{(l)}_{i}\cdot x^{(l)}_{i},\quad x^{(l+1)}=f^{(l)}(z^{(l)}),$
the corresponding trust computation is given by:
$T_{z^{(l)}}=\bigvee_{i=1}^{d}T_{x_{i}^{(l)}}\otimes T_{\theta_{i}^{(l)}},\quad T_{x^{(l+1)}}=T_{f}^{(l)}(T_{z^{(l)}}),$
where:
  • $\otimes$ is a trust discounting operator,
  • $\bigvee$ and $\oplus$ is a trust fusion operator,
  • $T_{f}^{(l)}$ is the trust-equivalent of the activation function. In this work, we set it to identity function as we use ReLU as activation function.
Motivated by the structure of trust propagation in earlier sub-problems (Eq. 14), the discount operator $\otimes$ models how trust in an input is modulated by trust in the associated parameter, while the fusion operator $\bigvee$ combines these trust contributions across inputs. The fusion operators should be associative or generalizable [43] to support multiple inputs.
With these definitions in place, we integrated Trust Nodes and Trust Functions into a parallel trust reasoning framework that mirrors neural network computation. Although this was first illustrated in the perceptron case, two critical challenges remain: how to quantify the trustworthiness of model parameters learned from datasets of variable quality, and how to generalize trust propagation to deeper, more complex neural architectures. These challenges motivate the design of the Parallel Trust Assessment System (PaTAS), a scalable architecture for trust propagation in neural networks. The following section presents its design, components, and theoretical foundations.
V PaTAS For Neural Networks
When neural network parameters such as $\theta_{1}$ and $\theta_{2}$ are learned (e.g., by backpropagation), their trustworthiness depends on the data used for training. If the dataset contains mislabeled or biased samples, parameter trust will be compromised. Yet this is only part of the problem. As discussed in Section II-A, Subjective Logic represents trust as a subjective binomial opinion with the three components trust (belief), distrust (disbelief), and uncertainty. This decomposition allows us to distinguish between different causes of unreliability: distrust may arise from systematic issues such as mislabeled or poisoned data, while uncertainty reflects variability or noise in the data. A central challenge, therefore, is how to exploit this advantage of subjective logic in order to make this distinction in practice.
V-A PaTAS General Description
To address the previous challenges, we introduce the Parallel Trust Assessment System (PaTAS), a framework designed to systematically propagate trust assessments through a neural network. PaTAS operates in parallel with the standard neural architecture and maintains a corresponding structure that mirrors the network’s topology. It enables the computation of trust in the output by integrating two key sources:
  1. 1.
    the trust in the input features at inference time, and
  2. 2.
    the trust in the parameters, derived from the training dataset trustworthiness calculated using a trust assessment function. This includes both the trust in the input features of the training samples and the trust in the corresponding labels.
To extend the trust-propagation principles introduced for the perceptron in the previous section, PaTAS organizes its Trust Nodes into a structure that mirrors the full neural network. Each neuron is associated with a corresponding Trust Node, and these nodes are connected according to the network’s topology. This generalization forms what we refer to as a Trust Nodes Network (TNN).
Definition 7 (Trust Nodes Network).
The Trust Nodes Network is a structured composition of Trust Nodes, where each Trust Node corresponds to a neuron in the underlying neural network. This network mirrors the architecture of the neural model and is responsible for propagating trust values across layers. The Trust Nodes Network computes trust assessments at each stage of inference by applying trust-specific reasoning operations aligned with the neural computation flow.
The PaTAS is designed to continuously evaluate the trustworthiness and preservation of properties, such as accuracy, during the inference of a neural network. For example, an input $x$ might be accurate, unbiased, and trustworthy (i.e., high trust score value $T_{x}$), while the corresponding output $y\prime$ may still be unreliable due to biased or poorly calibrated parameters $\theta$ (i.e., low trust score value $T_{y\prime}$). Rather than altering the neural network itself, PaTAS operates alongside it to evaluate the trustworthiness of computations (see Fig. 5). Embedding trust operations inside the network would modify its internal computations, risk harming accuracy, and substantially increase the computational cost. For this reason, PaTAS performs trust assessment in parallel, ensuring that trust reasoning does not interfere with the model’s predictions or complicate its operation.
Figure 5: Overview of PaTAS Workflow
V-B PaTAS Design
Figure 6: High-Level Overview of PaTAS Integration with a Neural Network
Figure 7: Functional flow of the PaTAS framework integrated with a neural network, illustrating how trust is propagated and revised during feedforward, backpropagation, and inference.
Feedforward (Training Phase): The NN interface extracts trust scores from input features for all data being processed. These trust scores (TF.1) are passed to the Trust Feedforward module, which propagates them through the Trust Nodes Network in parallel with the neural computations. During training, Trust Feedforward also stores the trust scores of all intermediate computations in the NN, to be reused in backpropagation.
Backpropagation (Training Phase): Once the NN computes gradients, the trust scores of the labels (B.1) and the gradients are provided to the Parameter-Trust Update. Gradients (B.2) indicate how dataset labels affect parameter updates, while label trust determines whether these updates should be considered reliable. The stored intermediate trust values are also incorporated in this process. Finally, the Parameter-Trust Update module refines the trust values of the Trust Nodes Network parameters (B.3), aligning them with the evolving learning dynamics.
Inference Phase: During operation, contextual information such as the set of activated neurons is retrieved. This information is passed to GenIPTA (I.1), which constructs an Inference Path Trust Assessment (IPTA) corresponding to the specific inference path (I.2). The IPTA takes input trust scores (I.3), propagates them along the path, and produces trust scores for each NN output. These can then be passed to the Output-Trust Aggregation (I.4) which will combine them into a single consolidated trust score representing the reliability of the prediction (I.5).
As illustrated in Fig. 6, PaTAS is composed of four main modules: Trust Feedforward, Output-Trust Aggregation, GenIPTA, and Parameter-Trust Update. Among these, the Parameter-Trust Update requires a more detailed treatment, since its design parallels the role of backpropagation in neural network training and involves elaborate reasoning operations. We therefore dedicate a separate subsection to it.
V-B1 Trust Feedforward
The Trust Feedforward function propagates trust values through the Trust Nodes Network in alignment with the neural network’s inference flow. It mirrors the layer-wise computations of the original model, but operates entirely on trust values. At each layer, trust in the inputs and parameters is combined using the trust operations defined in the Trust Function (including discounting and fusion), capturing how evidence flows through the network. This process produces a trust opinion for each output neuron of the neural network while also storing intermediate trust scores, which are later used by the Parameter-Trust Update.
V-B2 Output-Trust Aggregation
The Trust Feedforward function produces a vector of trust values, one for each output neuron. The goal of Output-Trust Aggregation is to combine these individual opinions into a single, consolidated trust score representing the overall trust in the network’s output. This combination is performed using a SL fusion operator, which fuses the trust opinions of all output Trust Nodes into one aggregated opinion. Alternatively, instead of aggregating across all outputs, one may directly consider the trust assigned to the final decision. For instance, in digit classification, if the model outputs a probability $0.9$ for class ‘1’, the trust in the decision can be taken as the trust score associated with the output neuron for class ‘1’.
V-B3 Inference-Path Trust Assessment Generation(GenIPTA)
The purpose of the GenIPTA module is to dynamically construct a function tailored to a single, specific inference. This function, called Inference-Path Trust Assessment (IPTA), reflects how trustworthy the exact computational path taken during that inference is. When an inference is performed, contextual information is recorded, such as the list of neurons activated along the path. In this case, the activation trace is used to instantiate a temporary subnetwork of Trust Nodes containing only the Trust Nodes corresponding to those activations, thereby mirroring the precise inference path of the neural network.
Contextual information is not limited to activation traces. Typically, all neurons are activated to some degree, but only a subset of these activations is strongly relevant to the decision. GenIPTA can therefore operate on truncated activation sets that retain only the most relevant neurons, which allows for more focused and interpretable trust assessments of the actual decision-making process. The GenIPTA can also work in a more complex way by using the actual activation values of all neurons, performing a weighted trust assessment where stronger activations contribute more heavily to the inference.
The detailed functional flow of PaTAS across feedforward, backpropagation, and inference is illustrated in Fig. 7.
V-C Parameter-Trust Update
Algorithm 1 Parameter-Trust Update Algorithm
1:Function ParameterTrustUpdate($g,T_{y},\epsilon$)
2:  Summary: Revises and updates the trust parameters of the Trust Nodes by combining gradient evidence, label trust, and neuron input trust, under mini-batch training.
3:  Step 1: Compute the aggregated trust in the labels
4:  $T_{y_{\text{batch}}}\longleftarrow\bigwedge_{y\in y_{\text{batch}}}T_{y}$
5:  for each layer $l$ do
6:   for each neuron $n_{i}^{(l)}$ do
7:     Step 2: Gather gradient evidence $g$ for neuron $n_{i}^{(l)}$
8:     $g^{(l)}_{i}\longleftarrow\{\,g^{(l)}_{i,j}\;|\;j\in\mathcal{N}(i)\,\}$
9:     Step 3: Compute trust for neuron $n_{i}^{(l)}$
[…]
11:     Step 4: Deduce overall trust in neuron $n_{i}^{(l)}$
12:     $T_{n_{i}\parallel Y_{\text{batch}}}\longleftarrow$ DeduceTrust($T_{n_{i}\mid y_{\text{batch}}},T_{n_{i}\mid\overline{y_{\text{batch}}}},T_{y_{\text{batch}}}$)
13:     for each incoming edge $j$ to $n_{i}^{(l)}$ do
14:      Step 5: Revise parameter trust with node trust
[…]
16:      Step 6: Update parameter trust with auxiliary factors
[…]
18:     end for
19:   end for
20:  end for
21:end Function
[…]
23:  Count $r$: #edges with $|g^{(l)}_{i,j}|<\epsilon$
24:  Count $s$: #edges with $|g^{(l)}_{i,j}|\geq\epsilon$
25:  Map $(r,s)$ into a binomial opinion using Baseline-Prior Quantification
26:  return $T_{n_{i}\mid y_{\text{batch}}}$
27:end Function
28:Function DeduceTrust($T_{n_{i}\mid y_{\text{batch}}},T_{n_{i}\mid\overline{y_{\text{batch}}}},T_{y_{\text{batch}}}$)
29:  $T_{n_{i}\mid\overline{y_{\text{batch}}}}\longleftarrow(0,0,1)$
30:  return $T_{n_{i}\parallel Y_{\text{batch}}}=T_{y_{\text{batch}}}\circledcirc(T_{n_{i}\mid y_{\text{batch}}},T_{n_{i}\mid\overline{y_{\text{batch}}}})$
31:end Function
[…]
33:  return $T_{\theta_{i,j}^{(l)}}\circleddash T_{n_{i}\parallel Y_{\text{batch}}}$
34:end Function
The notation used throughout this subsection and in the Algorithm 1 is summarized in Table VII in Appendix A.
The design of the Parameter-Trust Update is motivated by the standard backpropagation algorithm used in neural network training. In backpropagation, the gradient of the loss function with respect to the network parameters indicates how much each parameter contributes to the output error. Thus, for a parameter $\theta$, the update is given by
$\theta\leftarrow\theta-l_{r}\,\frac{\partial\mathcal{L}}{\partial\theta},$ (15)
where $l_{r}$ is the learning rate, $\mathcal{L}$ is the loss function and $\frac{\partial\mathcal{L}}{\partial\theta}$ is the gradient. The gradients are obtained recursively through the chain rule:
$\frac{\partial\mathcal{L}}{\partial\theta^{(l)}_{i,j}}=\delta^{(l)}_{i}\,x^{(l-1)}_{j}$, (16)
where $\theta^{(l)}_{i,j}$ denotes the weight connecting neuron $j$ in layer (l1)(l-1) to neuron $i$ in layer $l$, $\delta^{(l)}_{i}$ is the error term of neuron $i$ in layer $l$, and $x^{(l-1)}_{j}$ is the activation of neuron $j$ in the previous layer.
The Parameter-Trust Update mechanism integrates training signals, specifically the gradients $g$ computed during back propagation, together with trust in the labels in the training batch and the trust scores of intermediate computations stored during feedforward. The Parameter-Trust Update ensures that parameter trust reflects both observed model behavior and trust in the training data.
The Parameter-Trust Update process is formally described in Alg. 1. For each layer in the network, the framework revises the trust parameters of every Trust Node based on the evidence available from the training batch, the input processed during feedforward, and the gradients. It then updates these trust values to reflect the reliability of the newly updated parameters, using evidence from the learning rate and the intermediate trust scores.
The algorithm runs once per batch and begins by computing a combined trust opinion over all labels in the current batch (Line 3):
$T_{y_{\text{batch}}}=\bigwedge_{y\in y_{\text{batch}}}T_{y}$
This combined opinion serves as a foundation for conditioning the trust in each parameter.
For each neuron $n_{i}^{(l)}$ at index $i$ in layer $l$, the framework computes:
  • $T_{n_{i}^{(l)}\mid y_{\text{batch}}}$ (Line 9), the trust conditioned on the current batch labels. This is inferred by checking each incoming weight gradient $g^{(l)}_{i,j}$: if $|g^{(l)}_{i,j}|<\epsilon$ [^2] The threshold $\epsilon$ is not a tuned hyperparameter but a sensitivity parameter used only in the NodeTrust function to distinguish weak from strong gradients. Its value can be chosen in different reasonable ways depending on the desired sensitivity; in this work, we select $\epsilon$ relative to the typical gradient scale during training. Importantly, $\epsilon$ does not affect model predictions or training dynamics, but only the sensitivity of the trust-update mechanism. it is counted as positive evidence $r$, otherwise as negative evidence $s$. These counts are then mapped into a binomial opinion using the Baseline-Prior Quantification model.
  • $T_{n_{i}^{(l)}\mid\overline{y_{\text{batch}}}}$, the trust when the true batch labels are not $y_{\text{batch}}$. Since no concrete evidence is available for this case, it is initialized to a vacuous opinion: $(0,0,1)$.
[…]
Then for each incoming edge $j$ of neuron $i$, the parameter trust $T_{\theta_{i,j}^{(l)}}$ is updated in two stages:
  1. 1.
    Revision with deduced trust using fusion (Line 14):
    $T_{\theta_{i,j}^{(l)}}\leftarrow T_{\theta_{i,j}^{(l)}}\circleddash T_{n_{i}^{(l)}\parallel Y_{\text{batch}}}$
  2. 2.
    Adjustment with auxiliary factors (Line 16): During training, the update of parameter $\theta_{i,j}^{(l)}$ depends not only on its current value but also on auxiliary factors such as the learning rate $l_{r}$, the input feature $x^{(l-1)}_{j}$, and the label $y_{\text{batch}}$ (see Eqs. 15 and 16). To reflect this, we adjust parameter trust using:
    $T_{\theta_{i,j}^{(l)}}\longleftarrow T_{\theta_{i,j}^{(l)}}\odot(T_{x_{j}^{(l)}}\oslash T_{y_{\text{batch}}}).$
[…]
    $(b,d,u)$ $=(b_{1},d_{1},u_{1})\oslash(b_{2},d_{2},u_{2}),$ (17)
    $b$ $=\min(b_{1},b_{2}),$
    $d$ $=\max(d_{1},d_{2}),$
    $u$ $=1-(b+d).$
    This formulation reflects the fact that both unreliable features and mislabeled data can strongly bias parameter updates. The use of min and max provides a conservative aggregation: trust cannot exceed the weakest evidence, and distrust must reflect the strongest warning. It ensures that trust in both input features and labels is explicitly propagated into the parameter trust.
This process refines the trust in each parameter by incorporating both gradient-based behavioral evidence and auxiliary trust factors, allowing the PaTAS to align parameter trust with the training dynamics of the neural network.
V-D Theoretical Properties of PaTAS
To ensure that PaTAS provides reliable and interpretable trust assessments, we first establish several fundamental theoretical properties describing its stability and logical consistency.
V-D1 Convergence of PaTAS
The convergence of the PaTAS is governed by the stability of its inputs and the structure of its recursive Parameter-Trust Update process. During training, PaTAS updates the internal trust opinions associated with network parameters using trust opinions on the inputs, labels, hyperparameters, and gradient information. This update mechanism is outlined in Alg. 1. As proved in Theorem 1, the PaTAS converges under specific situation. This convergence relies on some specific characteristic of the operators used to feedforward and revise the trust in the parameters $\theta$.
Theorem 1 (Convergence of PaTAS Creation).
Let a neural network be trained until convergence, and let its associated PaTAS operate with:
  • a stable input trust assessment $T_{x}$,
  • a stable label trust assessment $T_{y}$,
  • a stable hyperparameter trust $T_{lr}$,
Let $T_{\theta}^{(n)}$ denote the trust opinion at iteration $n$ for parameter $\theta$. If the revision of the trust in the weights is performed as specified in Alg. 1, then the PaTAS parameter will also converge.
Proof.
See Appendix C
V-D2 Symmetry and Invariance Properties of PaTAS
Theorem 2 (PaTAS Inference on Vacuous Input Yields Vacuous Output).
Let $\omega^{\emptyset}=(0,0,1,a)$ denote a vacuous binomial opinion over any variable, with arbitrary base rate $a\in[0,1]$. Then the following two properties hold:
  1. 1.
    Discounting a Vacuous Opinion Yields a Vacuous Opinion.
    For any trust value binomial opinion $\omega^{A}_{B}$, the discounted opinion
    $\omega^{[A;B]}_{X}=\omega^{A}_{B}\otimes(0,0,1,a)=(0,0,1,a)$
  2. 2.
    PaTAS Feedforward on Vacuous Input Yields Vacuous Output.
    Let $T_{x}=(0,0,1,a)$ be the trust assessment of an input to PaTAS. Then for any parameter trust configuration $T_{\theta}$, the output trust assessment satisfies:
    $T_{y}=\text{IPTA}(T_{x})=(0,0,1,a).$
Proof.
See Appendix C
Definition 8 (Symmetric Binomial Opinions).
Let $x=(b,d,u,a)$ be a binomial opinion with belief $b$, disbelief $d$, uncertainty $u$, and base rate $a$ where $b+d+u=1$. The opinion $\bar{x}=(d,b,u,a)$ is called the symmetric of $x$. Two binomial opinions $x$ and $\bar{x}$ are symmetric if they share the same uncertainty and base rate and have inverted belief and disbelief, i.e.,
$x=(b,d,u,a),\quad\bar{x}=(d,b,u.a)\quad\text{with }b+d+u=1.$
Theorem 3 (Symmetric Inference under PaTAS).
Let $T_{\theta}$ be a fixed PaTAS inference operator based on subjective logic, and let $x=(b,d,u)$ be any binomial opinion with symmetric counterpart $\bar{x}=(d,b,u)$. Then the outputs $y=T_{\theta}(x)$ and $\bar{y}=T_{\theta}(\bar{x})$ are also symmetric, i.e.,
$y=(b^{\prime},d^{\prime},u^{\prime}),\quad\bar{y}=(d^{\prime},b^{\prime},u^{\prime}).$
In particular, the uncertainty is preserved:
$u_{y}=u_{\bar{y}},$
and the belief in one output equals the disbelief in the other:
$b_{y}=d_{\bar{y}},\quad d_{y}=b_{\bar{y}}.$
Moreover, for the fully trusted input $x=(1,0,0)$, the output satisfies $d_{y}=0$, and for the fully distrusted input $\bar{x}=(0,1,0)$, the output satisfies $b_{\bar{y}}=0$.
Proof.
See Appendix C
These symmetry and invariance properties serve as fundamental consistency checks, ensuring predictable behavior under neutral, vacuous, or balanced evidence, while also simplifying evaluation by reducing the number of distinct trust scenarios that need to be considered.
VI Evaluation and Results
The goal of our evaluation is to validate the PaTAS both theoretically and empirically. Specifically, we aim to demonstrate that PaTAS produces interpretable trust estimates that (i) converge during training under specific conditions, (ii) respect symmetry and invariance properties, and (iii) are able to provide interpretable assessments under realistic conditions such as noisy features, corrupted labels, or adversarial perturbations.
Our evaluation approach combines controlled synthetic degradations with real-world datasets. We systematically vary the trustworthiness of inputs ranging from fully trusted, fully uncertain, to fully distrusted, and observe how the created PaTAS propagates these trust assessments through the network. For each scenario, we track three complementary metrics: trust mass (belief), uncertainty mass, and distrust mass (disbelief). We also compare these values against standard model accuracy (after training) to understand how input trust affects output reliability.
We conduct three experiments of increasing complexity based on three different datasets:
  1. 1.
    Breast Cancer Classification, to assess behavior on a small, tabular medical dataset.
  2. 2.
    MNIST Digit Classification, to evaluate PaTAS across multiple neural architectures under controlled uncertainty.
  3. 3.
    Poisoned MNIST, to evaluate PaTAS and IPTA in the presence of adversarial corruption and data poisoning.
VI-A Experimental Setup
VI-A1 Experiment 1 - Breast Cancer Classification
We use the Breast Cancer Wisconsin (Diagnostic) Dataset [44], containing 569 samples with 30 numeric features (e.g., radius, area, symmetry) derived from breast mass fine needle aspirates. The neural network classifies the tumors into benign or malignant categories. The neural network architecture consists of 30 input neurons, 16 hidden neurons, and 2 output neurons, with ReLU activation in the hidden layer and Softmax in the output. The model is trained for 15 epochs with a batch size of 64 and a learning rate of 0.2, achieving 98% accuracy when the data are not modified.
For the evaluation, we degrade the training data in controlled ways and assign corresponding Subjective Logic trust assessments. We consider three extreme trust profiles (fully trusted $(1,0,0)$, fully distrusted $(0,1,0)$, and fully uncertain $(0,0,1)$) for both input features and label. These are combined (for inputs and label assessment) to form nine combinations. These dogmatic and vacuous opinions serve as canonical boundary cases in Subjective Logic: they express maximal trust, maximal distrust, and maximal uncertainty, respectively. We additionally include two intermediate scenarios, introduced below, to illustrate how PaTAS behaves under partial trust and partial distrust.
In practice, feature and label trust degradation may arise from poor-quality imaging, human annotation errors, or flaws in the preprocessing pipeline. In all experiments, we simulate such degradations using controlled perturbation functions. For fully uncertain feature opinion, we introduce additive uniform noise
$x^{\prime}=x+\mu,\quad\mu\sim U(-\eta,\eta),$
where
$\eta=0.3\times\max(\text{features}).$
Each feature is perturbed with probability $0.3$. After noise addition, if $x^{\prime}$ lies outside the valid range $\min(\text{features}),\max(\text{features})]$, it is clipped to the corresponding boundary:
$x^{\prime}=\begin{cases}\min(\text{features}),&x^{\prime}<\min(\text{features}),\\ \max(\text{features}),&x^{\prime}>\max(\text{features}),\\ x^{\prime},&\text{otherwise}.\end{cases}$
To model distrust, we generate corrupted inputs by sampling from a uniform distribution over the feature space:
$x^{\prime}\sim U(\min(\text{features},\max(\text{features}),$
Label degradation is modeled analogously: uncertainty is introduced via random label noise, while full distrust is represented by a complete replacement of the labels.
Finally, to complement the boundary cases, we evaluate two intermediate trust scenarios:
  1. i.
    a mild degradation, where features are perturbed with probability 0.15 and the trust assessment is set to $(0.25,0,0.75)$, and
  2. ii.
    the same scenario for a fully uncertain input features and fully uncertain labels, but where the assessment is set to $(0.25,0.25,0.5)$, reflecting partial distrust and trust, instead of a fully uncertain opinion (0,0,1).
VI-A2 Experiment 2 - MNIST
In this experiment, we evaluate the PaTAS framework on the MNIST dataset [45], which consists of 60,000 training images and 10,000 test images, each representing a digit from 0 to 9. Each image has 784 features (28x28 pixels). The neural network classifies these images into one of the 10 digit classes.
We test three neural network architectures:
  • Architecture 1: 784 input neurons, 5 hidden neurons, 10 output neurons (784-5-10).
  • Architecture 2: 784 input neurons, 10 hidden neurons, 10 output neurons (784-10-10).
  • Architecture 3: 784 input neurons, 20 hidden neurons, 10 output neurons (784-20-10).
Although these architectures are not state-of-the-art for MNIST, they are sufficient to evaluate the behavior of PaTAS across different model sizes and to demonstrate how trust propagates through the network. Each model uses the ReLU activation function for the hidden layer and Softmax for the output. Models are trained for 10 epochs with a batch size of 18 and a learning rate of 0.001, achieving test accuracies of 89%. We evaluate using fully uncertain Training data trust assessment functions for both input features and labels, corresponding to the case where we have no knowledge about the dataset.
VI-A3 Experiment 3 - Poisoned MNIST
In this experiment, we evaluate the PaTAS framework on a poisoned version of the MNIST dataset, where one third of the training images are corrupted: labels of digits 6 and 9 are flipped, and at the same time a visible patch of fixed size is added at the top-left corner of the corresponding images. This combination follows common practices in data poisoning and backdoor attack scenarios [46]. The remaining two thirds of the data remain clean. This setup allows us to examine how PaTAS responds to the simultaneous presence of corrupted labels and adversarial triggers that can undermine both model performance and trust.
We use Architecture 3 from the previous experiment, consisting of an input layer with 784 neurons, one hidden layer with 20 neurons, and an output layer with 10 neurons (784-20-10).
For the poisoned dataset, trust is assigned as follows:
  • Pixels corresponding to the patch are considered distrusted, while all other pixels are trusted.
  • Labels for patched images of digits 6 and 9 are distrusted, while others are trusted.
This setup helps the PaTAS framework focus on potentially corrupted areas, while trusting the unaffected parts of the data.
Implementation Details for All Experiments
During inference, multiplication operations in the feedforward phase are implemented using SL trust discounting, while addition operations employ a generalized SL averaging fusion operator to support summations over multiple inputs (see Definition 6). Trust revision uses the same averaging fusion. The PaTAS framework and Neural Network were implemented in Python 3.9 using NumPy. Two main modules were developed: PrimaryNN, which handles network structure, training, and inference, and PaTAS, which implements trust assessment and propagation functions defined in the operational flow (Fig. 7). These modules interact to dynamically evaluate and update trust during feedforward and backpropagation. The full implementation, including all modules, dependencies, and experiment scripts, is available as supplementary materials, with detailed instructions for reproducing all experiments.
VI-B Results and Analysis
The evaluation of the Parallel Trust Assessment System (PaTAS) is conducted during the training phase of the neural network. After each iteration of training (i.e., a complete feedforward and backpropagation cycle), we assume that an inference is performed, and we assess the trustworthiness of the corresponding output. This assessment is carried out under three input trust profiles:
  • Fully Trusted Input: where the trust in the data is considered trusted (row 1 in each figure).
  • Fully Uncertain Input: where the trust in the data is considered uncertain (row 2 in each figure).
  • Fully Distrusted Input: where the data is assumed to be unreliable (row 3 in each figure).
We focus on these three profiles because they represent the extreme and most informative boundary cases of input trust.
For each of these three input types, we track and plot the evolution of three key metrics over the course of training:
  • Trust Mass: representing the level of confidence in the output.
  • Uncertainty Mass: representing the uncertainty associated with the assessment.
  • Distrust Mass: representing the level of disbelief in the output.
As a result, for each evaluation, 9 distinct plots are generated, corresponding to the 3 input types (fully trusted, fully uncertain, and fully distrusted) and each of the three key metrics (trust, uncertainty, and distrust).
All the results are depicted in Figures in Sections D-A, D-B, D-C, D-D, D-E, D-F, D-G and D-H (Appendix D) and summarized in Tables II, II, III and IV. They confirm several theoretical properties of PaTAS proven in Section V-D:
X Trust Assessment Y Trust Assessment $\epsilon=0.01$ $\epsilon=0.1$ Train (%) Test (%)
fully distrusted fully distrusted 0 0 63 38
fully distrusted fully uncertain 0 0 53 62
fully distrusted fully trusted 0 0 63 63
fully uncertain fully distrusted 0 0 96 3
fully uncertain fully uncertain 0.28 0.31 53 46
fully uncertain fully trusted 0.30 0.34 97 96.49
fully trusted fully distrusted 0 0 98 0.88
fully trusted fully uncertain 0.27 0.33 52 89
fully trusted fully trusted 0.79 0.90 98 99
(0.25, 0, 0.75) (0.25, 0, 0.75) - 0.38 71 90
(0.25, 0.25, 0.5) (0.25, 0.25, 0.5) - 0.17 53 46
TABLE I: Summary of final trust mass and corresponding train/test accuracies (in %) for the Breast Cancer Classification Experiment.
Hidden Neurons Trust $t$ Train (%) Test (%)
5 0.28 31 58
10 0.30 37 71
20 0.32 39 76
TABLE II: Summary of final trust mass for the MNIST Classification Experiment.
  • Convergence of Trust Assessment: when accuracy converges, trust assessment converges as expected.
  • Inference on fully uncertain input yields fully uncertain output: a fully uncertain input always produces a fully uncertain output.
  • Symmetric Inference: when input trust assessments are symmetric, the inference results remain symmetric.
For detailed results, we primarily focus on the three plots for each experiment) corresponding to feedforward of fully trusted input. Processing uncertain inputs naturally yields uncertain outputs, while symmetry implies that trusted and distrusted cases mirror each other. We also note that when processing fully trusted inputs, the distrust mass generally remains close to zero. Since trust, uncertainty, and distrust sum to 1, analyzing only the trust mass is sufficient in such cases. Therefore we keep only the trust mass after the training in Tables II, II, III and IV .
Experiment 1 – Breast Cancer Classification
When both features and labels are clean, the model steadily improves and achieves high accuracy. Clean features with corrupted labels cause accuracy to collapse, while noisy labels yield only moderate and unstable learning. With corrupted features, accuracy remains low in all cases, even with clean labels. Noisy features with clean labels still permit relatively strong learning, but performance breaks down when labels are corrupted or noisy. Importantly, corrupted labels mislead the model: training accuracy appears high while test accuracy remains poor, showing that the model learns, but learns the wrong mapping. In contrast, noisy labels prevent learning altogether, regardless of whether features are clean or noisy. Corrupted features eliminate the ability to learn in any setting.
We evaluated the system for two values of $\epsilon$ (0.1 and 0.01), as summarized in Table II. We recall that $\epsilon$ is the threshold used by the NodeTrust function in Alg. 1.
  • For $\epsilon=0.01$: fully uncertain $X$ stabilizes trust mass around 0.28 for uncertain $Y$ and 0.30 for trusted $Y$. Fully trusted $X$ yields rapid increases, stabilizing at $\sim 0.27$ for uncertain $Y$ and $0.79$ for trusted $Y$.
  • For $\epsilon=0.1$: the system becomes less sensitive to gradients. When $X$ is uncertain, trust mass stabilizes around 0.31 for uncertain $Y$, and 0.34 for trusted $Y$. Fully trusted $X$ gives 0.33 for uncertain $Y$, and 0.90 for trusted $Y$. No matter $\epsilon$ value, when $X$ or $Y$ are distrusted the trust mass rapidly falls to 0.
Overall, there is a strong positive correlation between trust mass and test accuracy. When $\epsilon$ is smaller, the trust level of $Y$ (trusted, uncertain) has a weaker influence on the final trust mass. Notably, setting $X$ as fully uncertain while keeping $Y$ trusted yields a higher trust mass (0.30) than the opposite case, where $X$ is trusted but $Y$ uncertain (0.27). Furthermore, replacing a fully uncertain opinion $(0,0,1)$ assessment for X and Y with a mixed assessment $(0.25,0.25,0.5)$ reduces the trust mass from $0.31$ to $0.17$, showing that distrust degrades performance more strongly than trust enhances it.
Experiment 2 – MNIST Dataset
In this experiment, PaTAS produces ten output trust opinions, one for each class. However they are almost the same so we record only one in the table. As shown in Table II, when features and labels are noisy, accuracy improves as model size increases. However, training accuracy remains consistently lower than test accuracy, with the gap widening for larger networks. This indicates that the models learn useful representations but cannot fully fit noisy training data, while still generalizing well on clean test samples. Test accuracy is consistently greater than training accuracy across all architectures, suggesting training noise makes the model underestimate its learning progress. Larger architectures yield more stable trust assessments, likely due to smaller average gradient magnitudes.
Experiment 3 – Poisoned MNIST Dataset
As in the previous experiment, PaTAS has ten output trust opinions. These values are informative because the reliability of the inference path vary across labels, especially when some are poisoned and others are not. In Table III, we report the trust mass associated with label 3 (a clean class) and label 6 (a poisoned class; see Section VI-A3). Alongside overall train and test accuracy, we also report accuracy on clean test samples of digits 3 and 6, as well as accuracy on intentionally poisoned versions of those test samples.
The results suggest that PaTAS is able to reflect differences in reliability between the clean and poisoned classes, even under moderate corruption. Across patch sizes, the trust mass assigned to the clean label consistently remains higher than that assigned to the poisoned label, indicating a separation in how PaTAS evaluates the reliability of their respective inference paths. This separation depends strongly on the size of the patch used for poisoning:
  • $1\times 1$: minimal effect, where the trust mass of poisoned labels remains low, with only a small difference compared to clean labels.
  • $4\times 4$ and $20\times 20$: clear separation between clean and poisoned trust scores.
  • $27\times 27$: collapse of trust assessment, reflecting domination by the trigger.
In order to evaluate the IPTA, we also trained a binary neural network on the poisoned dataset with a patch size of $4\times 4$, achieving train/test accuracy of approximately $93\%$. Table IV reports both accuracy and the corresponding trust opinions for clean digits (3 and 6) and poisoned digit 6 datasets. Clean samples maintain high accuracy with balanced trust and uncertainty masses, with slightly better trust–uncertainty balance for digit 3. In contrast, poisoned samples show a drastic accuracy drop, lower trust, and higher uncertainty, alarming for untrustworthy predictions. Furthermore, when we explicitly distrust the patch pixels while trusting the remaining inputs, the resulting trust opinion becomes $(0.35,0.1,0.55)$. These results show that PaTAS provides interpretable warnings about poisoned outputs.
Patch size Trust for 3 Trust for 6 Train (%) Test (%) Clean 3 (%) Clean 6 (%) Poisoned 3 (%) Poisoned 6 (%)
$(1\times 1)$ 0.92 0.89 78 81 82 35 76 39
$(4\times 4)$ 0.90 0.83 76 75 88 17 49 21
$(20\times 20)$ 0.375 0.3 78 75 87 41 0 37
$(27\times 27)$ 0.05 0.03 80 80 88 68 0 0
TABLE III: Summary of result for the Poisoned MNIST Classification Experiment.
Accuracy (%) Trust Distrust Uncertainty
Clean 3 93.14 0.484 0.0 0.516
Clean 6 93.52 0.482 0.0 0.518
Poisoned 6 6.19 0.449 0.0 0.551
Poisoned 6 0.350 0.100 0.550
(patch distrusted)
TABLE IV: IPTA results for a binary neural network trained on poisoned datasets with patch size $4\times 4$. Results are obtained by feedforwarding a fully trusted opinion, except for the last row where patch pixels are explicitly distrusted.
VII Discussion
VII-A Discussion of PaTAS Findings
$g\to$ $T_{\theta_{i}\mid y_{\text{batch}}}\to$ $T_{y_{\text{batch}}}\to$ $T_{\theta_{i}\mid\mid{Y_{\text{batch}}}}\to$
0 Trusted Trusted Trusted
Uncertain Uncertain
DisTrusted Uncertain
$\infty$ DisTrusted Trusted DisTrusted
Uncertain (0,0.5,0.5)
DisTrusted Uncertain
TABLE V: Asymptotic behavior of $T_{\theta_{i}\mid\mid{Y_{\text{batch}}}}$
PaTAS is a framework for evaluating the runtime trustworthiness of neural network outputs. Beyond dynamic inference assessment, it can also estimate a static trust opinion of the network itself. The core principle is that a trustworthy model should preserve trust: a fully trusted input producing a highly trusted output indicates that the model does not erode trust during inference. Accordingly, the overall model trustworthiness can be defined as the trust score of the PaTAS feedforward function ($PaTAS_{FF}$) under a fully trusted input:
$T(\text{NN})=PaTAS_{FF}((1,0,0)),$
which quantifies how well trust is maintained throughout the network. Using this definition, we compared the Inference Path Trust Assessment (IPTA) for benign and adversarial inputs in Experiment VI-A3, where the adversarial case used a $4\times 4$ patch injection. Results show a degradation in trust for adversarial samples: benign inputs achieved $t=0.484,d=0,u=0.516$, while patched inputs dropped to $t=0.449,d=0,u=0.551$. These observations demonstrate the sensitivity of PaTAS to local perturbations, offering a quantitative indicator of reduced reliability.
Although trust mass and accuracy are often correlated, our results reveal meaningful divergences. In the breast cancer task (Table II), assigning a mixed trust profile $(0.25,0,0.75)$ to all features and labels yields a final trust mass of $0.38$, higher than the fully uncertain-features case, yet the latter achieves better test accuracy (96.49% vs. 90%). Thus, higher trust mass does not always imply superior predictive performance. A similar effect appears in the poisoned MNIST experiments (Table IV): the clean digit “6” attains slightly higher accuracy than “3” (93.52% vs. 93.14%), but IPTA assigns higher trust to “3” ($t=0.484$) than to “6” ($t=0.482$) which was trained with partially poisoned training data.
These differences reflect that PaTAS evaluates the reliability of the inference path, not just predictive accuracy. The digit “6” was a poisoned class during training, so the parameters involved in predicting “6” are less trusted than those for “3,” even though the test samples are clean.
Interpretation of trust values, like accuracy, depends on the application. In high-stakes settings, a trust score of $(0.4,0,0.6)$ may be inadequate, whereas in less critical domains it may suffice. PaTAS enables such contextual interpretation by providing a unified, interpretable metric that can be tracked over time, compared across models, or evaluated under varying conditions. Overall, PaTAS complements traditional metrics by signaling fragility even when accuracy appears high and by identifying stability where accuracy slightly decreases. This duality underscores its relevance in safety-critical contexts where accuracy alone can be misleading.
In practice, precise dataset trustworthiness estimates may be unavailable, especially for large training datasets where the provenance of individual samples is difficult to establish. PaTAS can still operate in such cases by initializing dataset trust to a fully uncertain (vacuous) opinion. While this limits the use of prior trust evidence during training, it preserves the ability to propagate trust assessments at inference time. For a specific query, the circumstances under which the input features were collected are often easier to assess, making input-level trust more practical to obtain than dataset-level trust. PaTAS can therefore support meaningful trust evaluation even when explicit dataset provenance is missing. Likewise, if trust in a particular input cannot be assessed, a fully uncertain opinion may be used as a fallback.
A key technical factor affecting PaTAS behavior is the parameter $\epsilon$, which controls the threshold used to classify gradients as positive or negative evidence during training. As shown in Table II, if $\epsilon$ is too small, PaTAS may misinterpret gradients as large even when the model performs well. Consequently, trusted labels may reduce parameter trust, resembling the case $g\to+\infty$ in Table V. Proper calibration of $\epsilon$ is therefore crucial, and can be tuned per layer or gradually reduced as training converges to increase sensitivity to smaller gradients.
Finally, trust quantification for $T_{\theta_{i}\mid y_{\text{batch}}}$ in the Trust Update Algorithm 1 relies on a simple gradient-counting procedure. Although computationally efficient, it fixes the uncertainty component of the resulting binomial opinion based on the number of input neurons to $i$. This fixed-uncertainty formulation may not be optimal in all cases, suggesting the need for adaptive quantification schemes. Similarly, $T_{\theta_{i}\mid\overline{y_{\text{batch}}}}$ is set to a fully vacuous opinion $(0,0,1)$, which, while consistent with the absence of evidence, may not fully leverage prior knowledge when available.
VII-B Trust Assessment and AI Security Threats
AI systems are exposed to diverse attacks targeting different stages of the lifecycle. We can distinguish six key phases: data collection, cleaning and labeling, dataset assembly, network design, model training, and deployment for inference.
A major threat is the data poisoning attack [46], where adversaries insert malicious or mislabeled samples into the training data to induce targeted misclassifications. PaTAS mitigates this threat by performing trust assessments at the feature or instance level during training. By computing dataset trustworthiness and propagating input trust through training, PaTAS can flag unreliable predictions during deployment.
Another vector is model stealing, in which adversaries issue numerous queries to reconstruct or approximate a deployed model, threatening intellectual property and enabling downstream attacks. By monitoring the trust of queries and corresponding predictions, PaTAS can detect anomalous or low-trust query sequences indicative of model extraction attempts.
A third class, membership inference attacks, aims to infer whether a specific data point was part of the training dataset, posing privacy risks. PaTAS supports mitigation by assessing output trust: consistently high confidence and low uncertainty may signal overfitting or memorization, while balanced trust indicates healthy generalization. Thus, output trust scores can help reveal potential information leakage.
These examples show that trust vulnerabilities can emerge throughout the AI pipeline. The effectiveness of PaTAS depends on the granularity of its assessments, but by enabling trust evaluation across multiple stages, it provides a flexible, context-aware defense mechanism that strengthens AI resilience against diverse adversarial threats.
VIII Conclusion
This paper presented the Parallel Trust Assessment System (PaTAS), a framework that models and propagates trust in neural networks through Subjective Logic. PaTAS introduces a parallel computational structure based on Trust Nodes and Trust Functions, enabling trust propagation alongside standard feedforward and backpropagation processes. The proposed Parameter Trust Update and Inference-Path Trust Assessment (IPTA) mechanisms jointly quantify how input quality, learned parameters, and activation paths contribute to the overall trustworthiness of model predictions.
Experimental evaluations across real-world and adversarial datasets demonstrate that PaTAS can identify trust degradation under data poisoning and adversarial patching while maintaining interpretability and stability. Results further show that trust scores derived from PaTAS capture reliability aspects not reflected by accuracy alone, offering a complementary view of model performance and robustness.
Beyond accuracy estimation, PaTAS provides a unified probabilistic foundation for quantifying model reliability across the AI pipeline, from dataset trust to inference-level decision confidence. Future work will extend PaTAS to large-scale and multimodal architectures, explore adaptive trust quantification schemes, and incorporate the role of parameter magnitudes in trust reasoning. In particular, we plan to investigate how weight values can inform the generation of the Trust Node network for individual inferences, allowing trust propagation to reflect not only network topology but also parameter influence. Overall, PaTAS establishes a systematic path toward trustworthy, transparent, and reliable neural learning systems.
Several directions remain open for further investigation. First, the effect of activation functions on trust propagation has not been studied in detail. In our experiments we used ReLU, which has convenient properties for trust reasoning. Other activation functions, for example sigmoid, tanh, or GELU, may influence how trust is transformed across layers. It is also necessary to explore the impact of various in-between layers in neural networks, such as convolutional or transformer layers, on the PaTAS’ trust evaluation. Understanding these effects is an important step toward applying PaTAS in a broader range of architectures.
Second, the computational cost of PaTAS deserves a more systematic evaluation. PaTAS operates in parallel with the neural network, and this parallelism can be exploited to optimize both inference and the construction of trust assessments. Benchmarking runtime and memory consumption on larger models will help identify opportunities for further optimization.
Third, a quantitative comparison with DeepTrust and related trust assessment frameworks is planned. This comparison will help clarify conceptual differences and highlight cases where PaTAS provides advantages in terms of interpretability or performance.
Appendix A Subjective Logic Operators and Notation for Parameter-Trust Update
TABLE VI: Summary of Subjective Logic Operators Used in This Work
Symbol Name Definition / Equation
$\odot$ Binomial multiplication [9]
$\otimes$ Trust discounting $\begin{aligned} \omega^{[A;B]}_{X}&=\omega^{A}_{B}\otimes\omega^{B}_{X}\\ (b,d,u)\otimes(b^{\prime},d^{\prime},u^{\prime})&=(Pb^{\prime},\;Pd^{\prime},\;1-P(b^{\prime}+d)\\ P&=b+au\end{aligned}$
$\oplus$ Averaging fusion [13]
[…]
TABLE VII: Symbols Used in the Parameter-Trust Update Subsection and Algorithm
Symbol Meaning / Description
Inputs to the Algorithm
$g$ Collection of gradients for all parameters in the batch
$T_{y}$ Trust opinion on label $y$
$\epsilon$ Gradient sensitivity threshold in NodeTrust
Batch and Layer Quantities
$T_{y_{\text{batch}}}$ Aggregated trust over labels in the current batch
$l$ Layer index
$n_{i}^{(l)}$ Neuron $i$ in layer $l$
$\mathcal{N}(i)$ Set of incoming edges to neuron $i$
Gradient Evidence
$g^{(l)}_{i,j}$ Gradient of loss w.r.t. $\theta^{(l)}_{i,j}$
$g^{(l)}_{i}$ Gradient vector for neuron $n_{i}^{(l)}$
$r$ Count of weak gradients: $|g^{(l)}_{i,j}|<\epsilon$
$s$ Count of strong gradients: $|g^{(l)}_{i,j}|\geq\epsilon$
Trust Values Computed in the Algorithm
$T_{n_{i}\mid y_{\text{batch}}}$ Trust in neuron $n_{i}$ conditioned on batch labels
$T_{n_{i}\mid\overline{y_{\text{batch}}}}$ Trust in neuron under incorrect labels (vacuous)
$T_{n_{i}\parallel Y_{\text{batch}}}$ Deduced trust in neuron $n_{i}$
Parameter Trust
$T_{\theta^{(l)}_{i,j}}$ Trust opinion on parameter $\theta^{(l)}_{i,j}$
$T_{lr}$ Trust in the learning rate
$T_{x^{(l-1)}_{j}}$ Trust in the input feature to parameter $\theta^{(l)}_{i,j}$
Operators Used
$\bigwedge$ Batch-wise fusion of trust opinions
[…]
$\odot$ SL binomial multiplication
[…]
Appendix B Symbols and Parameters Used in PaTAS
TABLE VIII: List of Symbols and Parameters Used in this work
Symbol Meaning / Description
Neural Network Quantities
$x$ Input feature vector
$y$ Ground-truth label
$y^{\prime}=f_{\Theta}(x)$ Output of the neural network
$\Theta=(W,b,f)$ Neural-network parameters (weights, biases, activations)
$W^{(l)}$ Weight matrix at layer $l$
$b^{(l)}$ Bias vector at layer $l$
$f^{(l)}(\cdot)$ Activation function at layer $l$
$z^{(l)}$ Pre-activation vector of layer $l$
$x^{(l)}$ Activation vector of layer $l$
$n_{i}^{(l)}$ Neuron $i$ in layer $l$
$\theta^{(l)}_{i,j}$ Parameter from neuron $j$ in layer l1l\!-\!1 to neuron $i$ in layer $l$
$\delta^{(l)}_{i}$ Backpropagated error signal of neuron $i$ in layer $l$
$x^{(l-1)}_{j}$ Activation of neuron $j$ in the previous layer
$\mathcal{N}(i)$ Set of incoming neighbors of neuron $i$
Gradients and Learning Dynamics
$g$ Collection of gradients for all network parameters
$g^{(l)}_{i,j}$ Gradient w.r.t. weight $\theta^{(l)}_{i,j}$
$g^{(l)}_{i}$ Vector of gradients for all incoming parameters of neuron $n_{i}^{(l)}$
$l_{r}$ (or $\mathrm{lr}$) Learning rate
$\mathcal{L}$ Loss function
Trust Assessments and Trust Nodes
$T(x)$ Trust assessment of input features
$T_{y}$ Trust opinion on label $y$
$T_{y_{\text{batch}}}$ Aggregated trust opinion over all labels in a batch
$T_{\theta^{(l)}_{i,j}}$ Trust opinion on parameter $\theta^{(l)}_{i,j}$
$T_{x^{(l)}_{i}}$ Trust opinion on activation $x^{(l)}_{i}$
$T_{x^{(l)}_{j}}$ Trust opinion of feature contributing to edge $(j\to i)$
$T_{lr}$ Trust opinion on the learning rate
$T_{n_{i}\mid y_{\text{batch}}}$ Trust in neuron $n_{i}$ conditioned on batch labels
$T_{n_{i}\mid\overline{y_{\text{batch}}}}$ Trust in neuron given incorrect labels (vacuous opinion)
$T_{n_{i}\parallel Y_{\text{batch}}}$ Deduced trust in neuron $n_{i}$ after combining conditional evidence
Subjective Logic Opinions and Evidence
$\omega=(b,d,u,a)$ SL binomial opinion: belief, disbelief, uncertainty, base rate
$r$ Positive evidence count (from weak gradients)
$s$ Negative evidence count (from strong gradients)
$\epsilon$ Gradient sensitivity threshold in NodeTrust
Trust Operators (SL Operators + PaTAS-specific)
$\oplus$ SL fusion operator (cumulative or averaging fusion)
$\otimes$ SL trust-discounting operator
$\ominus$ SL opinion-revision operator
$\odot$ SL binomial multiplication
[…]
$\bigwedge$ Batch-wise fusion of trust opinions
PaTAS System Components
$\mathrm{Pf}_{\Theta}$ Parallel Trust Function (PaTAS trust feedforward)
TNN Trust Nodes Network (parallel trust architecture)
GenIPTA Generator of the Inference-Path Trust Assessment
IPTA Inference-Path Trust Assessment for a single inference
Appendix C Theorems and Proofs (Section V-D)
C-A Key theorems
Theorem 4 (Convergence of Subjective Logic Arithmetic Sequence).
Let $(\Omega,\ominus)$ be a group. $\Omega$ is a set of opinions specified as $0,1]^{4}$ Let $\omega_{n}=(b_{n},d_{n},u_{n},a_{n})\in\Omega$ be a sequence defined recursively by the operator $\ominus$ as follows:
$\omega_{n+1}=\omega_{n}\ominus q,$
where $q\in\Omega$ and the operator $\ominus$ is a fusion operator.
Then the sequence $(\omega_{n})$ converges in $0,1]^{4}$ to a limit:
$\begin{cases}\omega_{0},&\text{if }q=(0,0,1,a_{0}),\\ q&\text{otherwise}.\end{cases}$
Proof.
Let us define a distance $d$ on the space $\Omega\subseteq[0,1]^{4}$, based on the Euclidean norm (2-norm).
Assume that the recursive update is expressed as:
$\omega_{n+1}=\omega_{n}\ominus q,q\in\Omega$
where $\ominus$ denotes a fusion operator (e.g., cumulative or averaging fusion) that combines opinions $\omega_{n}$ and $q$. By the properties of subjective logic fusion, the result of this operation satisfies one of the following:
  1. 1.
    $d(\omega_{n+1},q)<d(\omega_{n},q)$, i.e., the new opinion is strictly closer to $q$, or
  2. 2.
    $d(\omega_{n+1},q)=d(\omega_{n},q)$ and $\omega_{n+1}=\omega_{n}$, meaning the sequence has reached a fixed point.
This behavior reflects the nature of fusion operators, which are designed to generate an opinion that represents a consistent aggregation of the two inputs.
In case 1, the distance to $q$ strictly decreases at each step. Since the 2-norm is bounded in $0,1]^{4}$, the sequence $(\omega_{n})$ is contained in a compact space and forms a Cauchy sequence. Therefore, it converges to the unique fixed point $q$.
In case 2, where $\omega_{n+1}=\omega_{n}$, the sequence remains constant and equal to $\omega_{0}$. This occurs, for instance, when $q=(0,0,1,a_{0})$, representing full uncertainty. In that case, most of the fusion operators (almost all except averaging fusion) has no effect, and the sequence stays fixed.
Thus, in both cases, the sequence $(\omega_{n})$ converges. ∎
Theorem 5 (Convergence of Subjective Logic Geometric Sequence).
Let $(\Omega,\odot)$ be a group defined as in Theorem 4. Let $\omega_{n}=(b_{n},d_{n},u_{n},a_{n})\in\Omega$ be a sequence defined recursively by the operator $\odot$ as follows:
$\omega_{n+1}=\omega_{n}\odot q,$
where $q\in\Omega$ and the operator $\odot$ is the binomial multiplication defined by:
$\begin{cases}b_{x\odot y}=b_{x}b_{y}+\dfrac{(1-a_{x})a_{y}b_{x}u_{y}+a_{x}(1-a_{y})u_{x}b_{y}}{1-a_{x}a_{y}},\\ d_{x\odot y}=d_{x}+d_{y}-d_{x}d_{y},\\ u_{x\odot y}=u_{x}u_{y}+\dfrac{(1-a_{y})b_{x}u_{y}+(1-a_{x})u_{x}b_{y}}{1-a_{x}a_{y}},\\ a_{x\odot y}=a_{x}a_{y}.\end{cases}$
Then the sequence $(\omega_{n})$ converges in $0,1]^{4}$ if $a_{q}<1$. In particular:
  • $a_{n}=a_{0}a_{q}^{n}\to 0$ as $n\to\infty$,
  • $d_{n}\to 1$ if $d_{q}>0$, and $d_{n}=d_{0}$ if $d_{q}=0$,
  • $b_{n}\to 0$ if $p_{q}=b_{q}+a_{q}u_{q}<1$, and $b_{n}=b_{0}$ if $p_{q}=1$.
Proof.
We analyze each component of $\omega_{n}$ separately.
1. Convergence of $a_{n}$: By definition, $a_{n+1}=a_{n}a_{q}$. Since $a_{q}\in[0,1[$, this is a geometric sequence:
[…]
2. Convergence of $d_{n}$: The recurrence relation is:
$d_{n+1}=d_{n}+d_{q}-d_{n}d_{q}=d_{n}(1-d_{q})+d_{q}.$
This is a first-order linear recurrence. If $d_{q}>0$, the sequence is increasing and bounded above by 1. Therefore:
$\lim_{n\to\infty}d_{n}=1\,(\text{the fixed point})$
If $d_{q}=0$, then $d_{n+1}=d_{n}=d_{0}$ for all $n$.
3. Behavior of $b_{n}$: The update equation for $b_{n+1}$ is rational function involving $a_{n}$, $b_{n}$, and $u_{n}$. As $a_{n}\to 0$, the update expressions simplify:
$b_{n+1}\approx b_{n}b_{q}+a_{q}b_{n}u_{q}=b_{n}(b_{q}+a_{q}u_{q})=b_{n}p_{q}$
this converges since the projected probability $p_{q}\in[0,1]$
To conclude with $u_{n}$, since we have $u_{n}=1-(b_{n}+d_{n})$ it will also converge
C-B Proof for Theorem 1
Proof.
Assume that the neural network training process converges, implying that weight updates become increasingly small, and the back propagation gradients $g$ stabilize. This stability in $g$ reflects the fact that the model has reached a minimum or stable loss value.
Moreover, the stability of $T_{x}$, $T_{y}$, and $T_{l_{r}}$ implies that trust inputs to the update mechanism are stable. Consequently, each new trust update for $T_{\theta}$ is computed using consistent and bounded evidence, which over time results in the stabilization of the subjective opinions assigned to $T_{\theta}$.
In detail:
  • $T_{x}$ converging implies ($T_{y\prime}$ converging assuming the same $T_{\theta}$) implies convergence of $T_{y_{batch}}$.
  • $g$ converges so $T_{\theta\mid y_{\text{batch}}}$ will remain the same.
  • since $T_{\theta\mid\mid{Y_{\text{batch}}}}$ is a deterministic calculation from the convergent terms $T_{\theta\mid y_{\text{batch}}}$ and $T_{y_{batch}}$, it follows that $T_{\theta\mid\mid{Y_{\text{batch}}}}$ also converges
[…]
  • The function $f_{upd}$ in our implementation is based on $T_{\theta}$, $T_{l_{r}}$ and $T_{x}$} which all converge.
Therefore, assuming convergence of $T_{x}$, $T_{y}$, $T_{lr}$, and $g$, the trust values for all PaTAS parameters stabilize as training progresses, proving convergence of PaTAS creation. ∎
C-C Proof for Theorem 2
Let $\omega^{\emptyset}=(0,0,1,a)$ denote a vacuous binomial opinion over any variable, with arbitrary base rate $a\in[0,1]$. Then the following two properties hold:
  1. 1.
    Discounting a Vacuous Opinion Yields a Vacuous Opinion.
    For any trust value binomial opinion $\omega^{A}_{B}$, the discounted opinion
    $\omega^{[A;B]}_{X}=\omega^{A}_{B}\otimes(0,0,1,a)=(0,0,1,a)$
    Proof. Using the trust discounting operator from Subjective Logic, we set $P$ to the projected probability of $\omega^{A}_{B}$:
    $\omega^{[A;B]}_{X}=\begin{cases}b^{[A;B]}_{X}=P\cdot 0=0,\\ d^{[A;B]}_{X}=P\cdot 0=0,\\ u^{[A;B]}_{X}=1-b^{[A;B]}_{X}-d^{[A;B]}_{X}=1,\\ a^{[A;B]}_{X}=a.\end{cases}$
    Hence, $\omega^{[A;B]}=(0,0,1,a)$.
  2. 2.
    PaTAS Feedforward on Vacuous Input Yields Vacuous Output.
    Let $T_{x}=(0,0,1,a)$ be the trust assessment of an input to PaTAS. Then for any parameter trust configuration $T_{\theta}$, the output trust assessment satisfies:
    $T_{y}=\text{IPTA}(T_{x})=(0,0,1,a).$
    Proof. The PaTAS feedforward mechanism computes for each neuron:
    $T_{z}=\bigvee_{i}\left(T_{x_{i}}\otimes T_{\theta_{i}}\right).$
    Since each $T_{x_{i}}=(0,0,1,a)$, and using the result from part (1), we get:
    $T_{x_{i}}\otimes T_{\theta_{i}}=(0,0,1,a),\quad\forall i.$
    Then, by fusion of vacuous opinions:
    $T_{z}=\bigvee_{i}(0,0,1,a)=(0,0,1,a).$
    This holds recursively through all layers of PaTAS, including the output layer, hence:
    $T_{y}=(0,0,1,a).$
C-D Proof for Theorem 3
Proof.
We prove the theorem in two steps.
(1) Symmetry of the Discount Operator:
Let $\omega_{\theta}=(b_{\theta},d_{\theta},u_{\theta})$ be a binomial opinion representing the trust weight associated with a connection in the PaTAS. Let $P$ denote the projected probability of $\omega_{\theta}$, defined as:
$P=b_{\theta}+au_{\theta},$
where $a$ is the base rate (typically $a=0.5$ for binary domains).
Let $x=(b,d,u)$ be any binomial opinion and $\bar{x}=(d,b,u)$ its symmetric counterpart. Then the trust discounting operation yields:
$\omega_{\theta}\otimes x=(P\cdot b,P\cdot d,1-P\cdot(b+d)),$
$\omega_{\theta}\otimes\bar{x}=(P\cdot d,P\cdot b,1-P\cdot(b+d)).$
Since b+d=1ub+d=1-u, both discounted opinions have identical uncertainty and symmetric belief/disbelief masses. Hence, $\omega_{\theta}\otimes x$ and $\omega_{\theta}\otimes\bar{x}$ are symmetric.
(2) Symmetry Preservation under Fusion:
Let $\mathcal{X}=\{x_{1},\dots,x_{n}\}$ be a set of discounted opinions resulting from symmetric inputs, and let $\bar{\mathcal{X}}=\{\bar{x}_{1},\dots,\bar{x}_{n}\}$ be their symmetric counterparts. Consider any symmetric fusion operator $\bigoplus$ (such as averaging, cumulative fusion, or consensus fusion in Subjective Logic) applied to $\mathcal{X}$ and $\bar{\mathcal{X}}$.
Since each pair $(x_{i},\bar{x}_{i})$ is symmetric and the operator treats belief and disbelief symmetrically, we have:
$\bigoplus_{i=1}^{n}x_{i}=(b^{\prime},d^{\prime},u^{\prime})\quad\Rightarrow\quad\bigoplus_{i=1}^{n}\bar{x}_{i}=(d^{\prime},b^{\prime},u^{\prime}).$
Thus, the output of the PaTAS feedforward inference remains symmetric when symmetric inputs are provided.
(3) Invariance under Fusion:
A fundamental property of Subjective Logic fusion operators is that if all input opinions assign the same value to a specific mass (e.g., belief or disbelief), the result will preserve that value. In particular, if all input opinions have belief mass equal to zero, the fused result will also have belief mass zero. The same applies to disbelief mass.
Therefore, if the discounting step yields discounted opinions with zero belief (or zero disbelief) across all components, then the fusion stage will preserve that zero mass in the final output.
(4) Fully Trusted and Distrusted Cases:
For $x=(1,0,0)$ (fully trusted), we have:
$\omega_{\theta}\otimes x=(P,0,1-P),$
so the discounted opinion assigns zero disbelief. As noted above, the fusion of such discounted opinions will also assign zero disbelief.
For $\bar{x}=(0,1,0)$ (fully distrusted), we have:
$\omega_{\theta}\otimes\bar{x}=(0,P,1-P),$
so the discounted opinion assigns zero belief. Therefore, the belief from a fully distrusted input is always zero in the PaTAS output. ∎
Appendix D Detailed results
All the results are depicted in Figures in Sections D-A, D-B, D-C, D-D, D-E, D-F, D-G and D-H.
D-A Plots for Cancer Model with $\epsilon=0.01$ (Section VI-A1)
Figure 8: xdistrust+ydistrust
Figure 9: xdistrust+yvacuous
Figure 10: xdistrust+ytrust
Figure 11: xvacuous+ydistrust
Figure 12: xvacuous+yvacuous
Figure 13: xvacuous+ytrust
Figure 14: xtrust+ydistrust
Figure 15: xtrust+yvacuous
Figure 16: xtrust+ytrust
D-B Cancer Model with $\epsilon=0.1$(Section VI-A1)
Figure 17: xdistrust+ydistrust
Figure 18: xdistrust+yvacuous
Figure 19: xdistrust+ytrust
Figure 20: xvacuous+ydistrust
Figure 21: xvacuous+yvacuous
Figure 22: xvacuous+ytrust
Figure 23: xtrust+ydistrust
Figure 24: xtrust+yvacuous
Figure 25: xtrust+ytrust
D-C Results for trust Degradation (Section VI-A1)
Figure 26: (0.25, 0, 0.75)
Figure 27: (0.25, 0.25, 0.5)
D-D Accuracy Evolution for the Cancer Model (Section VI-A1)
(a) Clean Features and Labels
(b) Clean Features and Corrupted Labels
(c) Clean Features and Noisy Labels
(d) Noisy Features and Clean Labels
(e) Noisy Features and Corrupted Labels
(f) Noisy Features and Noisy Labels
(g) Corrupted Features and Clean Labels
(h) Corrupted Features and Corrupted Labels
(i) Corrupted Features and Noisy Labels
Figure 28: Accuracy evolution of the Cancer model under different combinations of clean, corrupted, and noisy features and labels.
Figure 29: Accuracy evolution of the Cancer model when reducing the noise for noised features and labels.
D-E MNIST with Vacuous Trust Assessment (Section VI-A2)
Figure 30: 5 Hidden neurons
Figure 31: 10 Hidden neurons
Figure 32: 20 Hidden neurons
Figure 33: 20 Hidden neurons with Fully Trusted Assessment
(a) 5 Hidden neurons
(b) 10 Hidden neurons
(c) 20 Hidden neurons
(d) 20 Hidden neurons with Fully Trusted Assessment
Figure 34: MNIST with Vacuous Trust Assessment
D-F MNIST poisoned 20 hidden neurons (Section VI-A3)
Figure 35: 1 pixel
Figure 36: 4×4 pixels
Figure 37: 20×20 pixels
Figure 38: 27×27 pixels
(a) 1 pixel
(b) 4×4 pixels
(c) 20×20 pixels
(d) 27×27 pixels
Figure 39: Accuracy for poisoned MNIST model with 20 hidden neurons
D-G MNIST poisoned 10 hidden neurons
Figure 40: 1 pixel
Figure 41: 4×4 pixels
Figure 42: 20×20 pixels
Figure 43: 27×27 pixels
D-H Example for random dataset trust assessment
Figure 44: MNIST randomized trust